OpenStudy (anonymous):

Help me plez? the line that passes through (-2,0) and has a slope of 5 in point slope form

OpenStudy (anonymous):

do you need a graph or the equation?

OpenStudy (anonymous):

the equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

woud it be y=8x+3??

OpenStudy (anonymous):

I mean y=-8x+3

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

so was it right or wrong?

OpenStudy (anonymous):

it was not correct

OpenStudy (anonymous):

whoops sorry that was my answer to the previous question

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

you should get something like y=5x+10

OpenStudy (anonymous):

so the answer for the line that passes through (-2,0) and has a slope of 5 in point slope form is y=5x+10?

OpenStudy (anonymous):

y-0=5(x+2) y-0=5x+10 y=5x+10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

another question: the line with a slope of -8 and a y intercept of 3 (in slope intercept form)

OpenStudy (anonymous):

would it come out to y=-8x+3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actually no

OpenStudy (anonymous):

okay thanks also How do I write a equation for: the horizontal line that passes through (-2,6)

OpenStudy (anonymous):

wait srry about that ignore the previous message

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the horizontal line equation is just y=6

OpenStudy (anonymous):

can u show how u got that? plez

OpenStudy (anonymous):

Whenever you want the equation of a horizontal line it is always the y value.

OpenStudy (anonymous):

same thing for the vertical line if you want the x value.

OpenStudy (anonymous):

so (-2,6) is in the form of (x,y)

OpenStudy (anonymous):

Ahh I see thanks

OpenStudy (anonymous):

Use the point-slope form linear equation: y+4=-4(x-2) Rewrite this equation in slope-intercept form.

OpenStudy (mayaal):

slope intercept form:y=mx+b m=slope (x,y)=point b=y-intercept

OpenStudy (mayaal):

just solve the above equation for y

OpenStudy (mayaal):

y=-4x+4

OpenStudy (mayaal):

did u get it?

OpenStudy (anonymous):

ur a life saver.. how do I write it in standerd form

OpenStudy (mayaal):

this is the standard form

OpenStudy (anonymous):

then how do I write it in slope intercept form?

OpenStudy (mayaal):

y=-4x+4

OpenStudy (anonymous):

so its written the same way both times?

OpenStudy (mayaal):

yes.

OpenStudy (anonymous):

thanks a lot. what is the x-intercept of this line?

OpenStudy (mayaal):

ur wm:)u plug in 0 for y and then solve for x

OpenStudy (anonymous):

so it would be (0,-4)?

OpenStudy (mayaal):

no,its (0,1)

OpenStudy (anonymous):

can u show me how u got that?

OpenStudy (mayaal):

y=-4x+4 0=-4x+4 -4=-4x -1=x

OpenStudy (anonymous):

for the y intercept I got (0,8) is that right?

OpenStudy (mayaal):

no,its (0,4),and sorry,the x-intercept is(1,0).

OpenStudy (anonymous):

so at the end it would be 1=x?

OpenStudy (mayaal):

yes.u r correct:)

OpenStudy (anonymous):

lol ur the best! Im almost done. What is the equation in standard for of a perpendicular line that passes through(5,1)?

OpenStudy (anonymous):

are u still there?

OpenStudy (mayaal):

yes,i have to submit an assignment by 5pm so im trying to do my last question.sorry for the late reply:)

OpenStudy (anonymous):

its okay go ahead and finish ur assignment

OpenStudy (mayaal):

i don't understand ur question

OpenStudy (anonymous):

Use the point-slope form linear equation: y+4=-4(x-2) What is the equation in standard for of a perpendicular line that passes through(5,1)?

OpenStudy (mayaal):

y=1/4x-1/4 is the perpendicular line's equaton in slope intercept form

OpenStudy (anonymous):

:0 I have to write it in standard form.. and show work dude ur seriously the best :)

OpenStudy (mayaal):

y+4=-4(x-2) the slope of this line is -4,so the slope of the perpendicular line is 1/4. y-y1=m(x-x1) y-1=1/4(x-5) y-1=1/4x-5/4 y=1/4x-5/4+1 y=1/4x-1/4

OpenStudy (mayaal):

r u still confused?

OpenStudy (anonymous):

nope u really helped :DD

OpenStudy (anonymous):

what is the x intercept of the perpendicular line?

OpenStudy (mayaal):

(1,0)

OpenStudy (anonymous):

can u plez show me how u got that mayaal?

OpenStudy (mayaal):

sure.i substituted 0 for y,and then solved the equation to find x.

OpenStudy (anonymous):

I wish I can u a medal.. what is the equation in standard form of a parallel line that passes through (0,2)